• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name mitchell

    Log entry time 15:59:22 on June16,2002

    Entry number 86165

    keyword=shift summary

    8:05 Arrive to find Q2d_d40 run 1537 in progress

    run 1537 complete Q2d_d40
    run 1538 complete Q2d_d40
    run 1359 complete Q2d_d40 -- during run 1539 Left Q1 has helium level fault.

    8:44 Left Q1 helium fault
    8:54 Call Ed Folts: ask MCC for access
    9:15 Left Q1 Reset
    9:40 Talk to MCC, busy with injector clean-up: no time estimate
    10:01 Beam is back!!!
    10:02 During long down we changed to Q2d_d50 (including trigger setup).
    10:03 MCC calls and says that they have a vacumm problem: ETA ~30 min
    10:06 Left Q1 Helium Fault
    10:15 Q1 Reset
    10:45 Beam is back!!!
    10:47 Start run 1541 First run of Q2d_d50
    10:58 Stop run--Roc14 dies. Reboot from FUI works.
    11:03 Start run 1542 (production Q2d_d50)
    11:16 Stop run 1542
    11:17 setup for BCM calibration (cleared by Hall B)
    11:28 start BCM--run 1543
    12:38 BCM calibration complete !!!
    12:39 start run 1544 (production Q2d_d50)
    13:41 start run 1545 (production Q2d_d50)
    13:42 beam off -- 13:48 beam on
    14:48 start run 1546 (production Q2d_d50)
    15:48 start run 1547 (production Q2d_d50)